OpenStudy (dtan5457):

Solve for when 0<=x<2pi 2cos(x/2)+1=0

OpenStudy (dtan5457):

@dan815

OpenStudy (dtan5457):

@Nnesha

OpenStudy (dtan5457):

So far I'm getting just 240 degrees...what else..

OpenStudy (dtan5457):

oh, and 120, is that all?

OpenStudy (amistre64):

cos(x/2) = -1/2 x/2 = 180 +- 60 x = 360 +- 120

OpenStudy (amistre64):

convert to radians tho